40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 9 23:17:47 2010, kiwamu, Update, SUS, beam pointing has been done 
    Reply  Fri Dec 10 09:58:57 2010, Aidan, Update, SUS, beam pointing has been done 
       Reply  Sat Dec 11 00:41:55 2010, kiwamu, Update, IOO, alignment of IP_ANG is done 
          Reply  Tue Dec 14 01:04:23 2010, kiwamu, Update, SUS, alignment of ITMs and PRM done 
             Reply  Tue Dec 14 15:15:19 2010, kiwamu, Update, SUS, put ETMY and aligned the beam 
                Reply  Tue Dec 14 22:42:56 2010, kiwamu, Update, SUS, installed the last TT . And some more progresses 
                   Reply  Thu Dec 16 01:23:43 2010, kiwamu, Update, SUS, SRM OSEMs installed DSC_2757_ss.jpg
Message ID: 4039     Entry time: Thu Dec 9 23:17:47 2010     Reply to this: 4040
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: beam pointing has been done 

[Koji, Osamu and Kiwamu]

 We aligned the beam axis pointing down to both X and Y arm.

Now the beams are hitting the centers of both ETMX and ETMY.

Amazingly Osamu made X arm flashing by aligning the cavity.


(what we did)

- opened almost all the chambers except for the MC2 chamber.

- locked and aligned the MC.

   We set Marconi to the right frequency, which had been set to the default values, probably due to the power outage in the last weekend.

   Also we found a DAC cable disconnected from the IO chassis of c1sus. So we connected it in order to damp the MC suspensions.

- aligned MMT2 and PZT2 in order to let the beam go through the center of PRM.

- checked the beam centering at the two TTs (PR2, PR3).

- rotated PR3 to make the beam go through the centers of both ITMY and BS at the same time.

- tried finding the beam spot at the ETMY chamber, and successfully found it.

    To see such faint beam spot, we used an IR viewer.

    In addition to that, we put a large piece of aluminum foil as a screen in the chamber.

- aligned the beam to the center of ETMY by tweaking the PZT mirror (SM2).

- aligned the BS so that the reflected beam at the BS goes through the center of ITMX.

- tried finding the beam spot at the X end, and successfully found it hitting the wall in the chamber.

- aligned the BS in order to let the beam hit the center of ETMX.

- tried aligning ETMX and ITMX to the beam.

Eventually we made the X arm flashing.

However the flash was a bit too weak to completely align the cavity.

 

(plan for tomorrow)

- reinstall some steering mirrors into the BS chamber

- check and neutralize PZT1

- alignment of IP_ANG

 

ELOG V3.1.3-